# Unitary discriminants of characters

### v1

@@
Abstract. Together with David Schlang we computed the discriminants of the invariant Hermitian forms for all indicator $o$ even degree absolutely irreducible characters of the ATLAS groups supplementing the tables of orthogonal determinants computed in collaboration with Richard Parker, Tobias Braun and Thomas Breuer. The methods that are used in the unitary case are described in this paper. A character has a well defined unitary discriminant, if and only if it is unitary stable, i.e. all irreducible unitary constituents have even degree. Computations for large degree characters are only possible because of a new method called unitary condensation. A suitable automorphism helps to single out a square class of the real subfield of the character field consisting of representatives of the discriminant of the invariant Hermitian forms. This square class can then be determined modulo enough primes.
@@
keywords: unitary representations of finite groups; unitary discriminants; invariants of Hermitian forms; Clifford invariants of quadratic forms
@@ 1 Introduction
Let $K$ be a field, $G$ a finite group and $n\in{\mathbb{N}}$. A group homomorphism $\rho:G\to\operatorname{GL}_{n}(K)$ is called a $K$-representation of $G$. Often $\rho(G)$ is contained in a smaller classical group, such as symplectic, unitary or orthogonal groups. This paper continues a long term project of the author with Richard Parker to specify these classical groups for finite fields and number fields. The fundamental methods for studying orthogonal representations are described in our joint paper [13]. Parker was mainly interested in finite fields $K$. For given even dimension $n=2m$ there is a unique unitary and symplectic group over a finite field; however there are two non-isomorphic orthogonal groups, $O_{2m}^{+}(K)$ and $O_{2m}^{-}(K)$. So we were faced with the task: given a Brauer character $\chi$ of an irreducible orthogonal representation $\rho:G\to\operatorname{GL}_{2m}(K)$, decide whether $\rho(G)$ is contained in an orthogonal group $O_{2m}^{+}(K)$ or $O_{2m}^{-}(K)$, i.e. whether $\chi$ has orthogonal discriminant $O+$ or $O-$. The finitely many primes $p$ dividing the group order $|G|$ can be handled case by case, the infinitely many other primes $p$ are dealt with by computing the orthogonal discriminant of the corresponding indicator $+$ ordinary character $\chi$. This is possible because, for a given complex irreducible character $\chi$ of indicator $+$ and even degree, there is a square class of the character field of $\chi$ containing determinants of the $\rho(G)$-invariant quadratic forms, though, due to Schur indices, there might not be a representation $\rho$ over the character field that affords the character $\chi$ (see [12]). Together with Breuer and Parker, the author compiled a list of these orthogonal discriminants for all indicator + even degree ordinary and Brauer characters for all groups in the ATLAS of finite groups of order smaller or equal to the order of the sporadic simple Harada-Norton group. The results are available in [1]. As the computations seem to be much more involved, we left out to determine the Clifford invariant for the indicator + characters over number fields and also did not handle unitary discriminants for ordinary complex irreducible characters of indicator $o$ and even degree.
@@
The present paper transfers the ideas of [13] to the unitary case. Many methods, such as restriction, induction, tensor products and symmetrizations, behave very similar to the ones for orthogonal characters. Also decomposition number techniques can be varied to obtain restrictions for the unitary discriminants (see Section 5.2), for instance showing that all primes dividing the unitary discriminant also do divide the group order (see Corollary 5.5).
@@
However there is a great difference to the orthogonal case: The discriminant of a given $L/K$-Hermitian form $H$ is only well defined modulo norms, $\operatorname{disc}(H)\in K^{\times}/N_{L/K}(L^{\times})$. For finite fields, the norm map is surjective, so it seems to be impossible to obtain any information about $\operatorname{disc}(H)$ by reducing the representation modulo primes. The latter technique is important in [13] to handle large degree characters, for which we computed composition factors of a condensed module (see Section 8.1) modulo some well chosen primes (usually not dividing the group order) to get enough information on the square class containing the orthogonal discriminant. It is amazing that also this condensation method can be transferred to the unitary case: If there exists a suitable automorphism $\alpha$ of order 2 of $G$, interchanging the indicator $o$ character $\chi$ with its complex conjugate $\overline{\chi}=\chi\circ\alpha$, then there is a square class of the real subfield of the character field of $\chi$, called the $\alpha$-discriminant of $\chi$, that relates to the unitary discriminant of $\chi$ (see Theorem 7.6). In complete analogy to the orthogonal case, we can use skew adjoint units in the $\alpha$-fixed algebra, to compute the $\alpha$-discriminant of $\chi$ (see Section 8.2).
@@
We start by recalling the basic notions for quadratic and Hermitian forms. As we are mainly interested in totally positive definite Hermitian forms over quadratic complex extensions $L$ of totally real number fields $K$, Section 3 gives the important facts for this situation. In particular Hasse’s norm theorem gives criteria for computing the norm subgroup $N_{L/K}(L^{\times})$ of $K^{\times}$ from local data. Invariant lattices provide a tool to use modular reduction to restrict the possible unitary discriminants. Therefore some facts about lattices in local Hermitian spaces are recalled in Section 3.2. The next section introduces unitary discriminants of characters. To appropriately use modular reduction techniques, we need to have a finer notion of indicator for absolutely irreducible Brauer characters: There are two different sorts of indicator $o$ Brauer characters; those, for which the corresponding simple module carries a non-zero invariant Hermitian form, and those for which there is no such invariant form. The former characters are called unitary, as the corresponding representation embeds into the unitary group. To simplify notation we also say that all real Brauer characters and all ordinary characters are unitary. Then a character is called unitary stable (see Definition 5.1) if and only if all its unitary constituents have even degree. Unitary stable characters are exactly those that have a well defined unitary discriminant. In particular indicator -- characters are unitary stable and Section 4.3 shows how to determine their unitary discriminant from their local Schur indices. Section 5 collects important character theoretic methods that can be automatised to compute and check unitary discriminants. For orthogonal characters there was no need to deal with covering groups, as there the orthogonal discriminants are predicted by theoretical methods. For unitary characters the $Q_{8}$-trick described in Section 6 is a shortcut to obtain unitary discriminants of faithful characters of certain even covering groups.
@@
The next section lays the ground for applying condensation techniques to compute unitary discriminants. The determinant of a symmetric bilinear form can be computed as the determinant of its adjoint involution, and hence as the square class of the determinant of any skew-adjoint unit. The adjoint involution of an Hermitian form $H$ is an involution of second kind. Also here we can obtain the determinant of $H$ intrinsically from the algebra $\operatorname{End}(H)$ with involution (see Remark 4.8). There are two obstacles, however: The discriminant algebra from Remark 4.8 is hard to determine in general, and, the determinant of $H$ is only defined modulo norms. We overcome these obstacles by defining an orthogonal subalgebra (Definition 7.4), $A\leq\operatorname{End}(H)$, to which the adjoint involution of $H$ restricts as an orthogonal involution. Orthogonal subalgebras can be determined as $\alpha$-fixed algebras (Definition 7.2) for certain outer automorphisms $\alpha$. Section 8 shows how this can be applied to compute unitary discriminants using modular condensation techniques. This is illustrated in two examples, a computational one for the Harada-Norton group and a smaller example for $U_{3}(7)$, that illustrates techniques that might generalise to compute all unitary discriminants of the groups $U_{3}(q)$ for odd prime powers $q$ and other finite groups of Lie type. The last section illustrates the computation for the group $3.ON$, the Schur cover of the sporadic simple O’Nan group. Besides giving examples for the various methods described in this paper, the results are interesting, as this is one of the rare cases, where some unitary discriminants are not rational.
@@
Special thanks go to David Schlang, who compiled a large list of unitary discriminants during his master thesis [18] and his fellowship awarded by the SFB TRR 195, Computeralgebra (Project-ID 286237555). With his help we obtained the unitary discriminants of all indicator $o$ even degree ordinary characters for all groups in the ATLAS of finite groups [5] of order smaller or equal to the order of the Harada Norton group, the current status of the database of orthogonal discriminants. I also thank Thomas Breuer, who is permanently updating the OSCAR database of orthogonal and unitary discriminants of characters [1].
@@ 2 Quadratic and Hermitian forms
[…]
@@
For quadratic forms $H$, i.e. $L=K$, the square class $\operatorname{disc}(H)=d(K^{\times})^{2}$ defines a unique étale $K$-algebra $\mathcal{D}(H)=K[X]/(X^{2}-d)$ of degree 2 over $K$, the discriminant algebra of the quadratic form (see [13, Definition 3.1, Remark 3.2] also for the correct definition for fields of characteristic $2$). For $L\neq K$ the class of the discriminant defines a unique quaternion algebra over $K$ which we call the discriminant algebra of $H$:
@@
For $a,b\in K^{\times}$ let $(a,b)_{K}$ denote the central simple $K$-algebra with $K$-basis $(1,i,j,ij)$ such that $i^{2}=a$, $j^{2}=b$, $ij=-ji$. If $L=K[\sqrt{\delta}]$ is a quadratic extension of $K$ we also put
@@
where $\operatorname{Br}_{2}(L,K)$ is the exponent $2$-subgroup of the Brauer group of central simple $K$-algebras that are split by $L$.
@@
For $[D]\in\operatorname{Br}_{2}(L,K)$ the $L$-discriminant of $[D]$ is defined as
@@
where $b\in K^{\times}$ is such that $[D]=[(L,b)_{K}]$.
For an $L/K$-Hermitian form the $L$-discriminant of $\Delta(H)$ is just the discriminant of $H$.
@@ 3 Hermitian forms over number fields
[…]
@@
The following proposition is given in [17] for general global fields. Note that in [17, Remark (10.1.4)] the determinant of the Hermitian form has to be replaced by its discriminant to obtain a correct statement.
@@
(see [17, Remark (10.1.4), Theorem (10.1.7)]) Two Hermitian spaces $(V,H)$ and $(V^{\prime},H^{\prime})$ are isometric if and only if the quadratic spaces $(V_{K},Q_{H})$ and $(V^{\prime}_{K},Q_{H^{\prime}})$ are isometric. If $L=K[\sqrt{\delta}]$ then $\operatorname{disc}(Q_{H})=\delta^{n}$ and the Clifford invariant of $Q_{H}$ is $\Delta(H)$.
@@
Two totally positive definite $L/K$ Hermitian forms are isometric, if and only if they have the same dimension and the same discriminant algebra.
An important feature for number fields is the so called Hasse principle or local-global principle:
@@
  • Two quaternion algebras over $K$ are isomorphic, if and only if they ramify at the same places of $K$. Moreover the number of ramified places is always even.
  • Hasse Norm Theorem for quadratic extensions: For $a\in K$ it holds that $a\in N_{L/K}(L)$ if and only if $a\in N_{L_{\wp}/K_{\wp}}(L_{\wp})$ for all places Weierstrass-$\wp$ of $K$.
The finite completions of $K$ are in bijection to the maximal ideals of its ring of integers ${\mathbb{Z}}_{K}$. Let $\wp\unlhd{\mathbb{Z}}_{K}$ be such a non-zero prime ideal. Then there are three cases:
@@
    $\wp{\mathbb{Z}}_{L}$ is a maximal ideal of ${\mathbb{Z}}_{L}$, i.e. Weierstrass-$\wp$ is inert in $L/K$ and $L_{\wp}$ is the unique unramified quadratic extension of $K_{\wp}$.
@@
    $\wp{\mathbb{Z}}_{L}=P_{1}P_{2}$ is a product of two distinct prime ideals $P_{1}=\sigma(P_{2})$ of ${\mathbb{Z}}_{L}$, i.e. Weierstrass-$\wp$ is split in $L/K$ and $L_{\wp}\cong K_{\wp}\oplus K_{\wp}$.
@@
    $\wp{\mathbb{Z}}_{L}=P^{2}$ is the square of a prime ideal of ${\mathbb{Z}}_{L}$. Then Weierstrass-$\wp$ is ramified in $L/K$.
@@
    $\nu_{\wp}(a)$ is even for all prime ideals Weierstrass-$\wp$ of $K$ that are inert in $L/K$
@@
The first statement guarantees that $a$ is a norm for all infinite places of $K$, the second one that $a$ is a local norm for all inert places and the third one for the ramified places. For the split places $L_{\wp}=K_{\wp}\oplus K_{\wp}$ and all elements of $K_{\wp}$ are norms. ∎
For the infinite completions the complex signature parametrizes Hermitian forms. As $H$ is totally positive definite, this specifies $H$ for the infinite completions of $K$.
@@
The infinite places of $K$ are ramified in the discriminant algebra $\Delta(H)$ if and only if $\operatorname{disc}(H)$ consists of totally negative elements, if and only if $\dim_{L}(V)\equiv$ $2$ or $3$ mod $4$.
Let Weierstrass-$\wp$ be a prime ideal of $K$ that is split in $L/K$. Then the completion $\Delta(H)_{\wp}$ contains the completion $L_{\wp}\cong K_{\wp}\oplus K_{\wp}$ which contains zero divisors. So Weierstrass-$\wp$ is not ramified in $\Delta(H)$.
@@
Let Weierstrass-$\wp$ be a prime ideal of $K$ that is inert in $L/K$. Then Weierstrass-$\wp$ is ramified in $\Delta(H)$ if and only if the Weierstrass-$\wp$-adic valuation $\nu_{\wp}(d)$ is odd for all $d\in\operatorname{disc}(H)$.
@@
Now Weierstrass-$\wp$ is ramified in $\Delta(H)=(L,d)_{K}$ if and only if $d$ is not a norm in $L_{\wp}/K_{\wp}$, i.e. if and only if $\nu_{\wp}(d)$ is odd. ∎
@@
For a ramified place $\wp=P^{2}$ we can always find a representative $d\in{\mathbb{Z}}_{K}$ of $\operatorname{disc}(H)$ such that $d\not\in\wp$ (see Proposition 3.10 (c)). Assume that Weierstrass-$\wp$ is not a dyadic prime, i.e. $2\not\in\wp$. Then Weierstrass-$\wp$ ramifies in the discriminant algebra of $H$ if and only if $d$ is not a square modulo Weierstrass-$\wp$.

### v2

@@
Abstract. Together with David Schlang we computed the discriminants of the invariant Hermitian forms for all indicator $o$ even degree irreducible characters of ATLAS groups supplementing the tables of orthogonal determinants computed in collaboration with Richard Parker, Tobias Braun and Thomas Breuer. The methods that are used in the unitary case are described in this paper. An ordinary character has a well defined unitary discriminant algebra, if and only if it is unitary stable, i.e. all irreducible unitary constituents have even degree. Computations for large degree characters are only possible because of a new method called unitary condensation. Often, a suitable automorphism helps to single out a square class of the real subfield of the character field consisting of representatives of the discriminant of the invariant Hermitian forms. This square class can then be determined modulo enough primes.
@@
keywords: unitary representations of finite groups; unitary discriminants; invariants of Hermitian forms; Clifford invariants of quadratic forms; discriminant algebras of involutions of the second kind
@@ 1 Introduction
Let $K$ be a field, $G$ a finite group and $n\in{\mathbb{N}}$. A group homomorphism $\rho:G\to\operatorname{GL}_{n}(K)$ is called a $K$-representation of $G$. Often $\rho(G)$ is contained in a smaller classical group, such as symplectic, unitary or orthogonal groups. This paper continues a long term project of the author with Richard Parker to specify these classical groups for finite fields and number fields. The orthogonal groups are dealt with in the joint paper [15] and symplectic groups are essentially uniquely determined by the irreducible character. This paper approaches computational challenges to precise unitary representations over number fields. Whereas over finite fields and also over the real numbers there is a unique compact unitary group of any given dimension, there are infinitely many such groups over number fields if the dimension of the underlying Hermitian space is even. These unitary groups are uniquely determined by the discriminant of the invariant definite Hermitian forms.
To be more precise let $\chi$ be an ordinary indicator $o$ irreducible character of the finite group $G$ and put $L:={\mathbb{Q}}(\chi)$ to denote its character field, i.e. the number field that is generated by the character values of $\chi$. Then $L$ is a totally complex number field. Let $K$ denote its real subfield and $N:=N_{L/K}(L^{\times})$ the norm subgroup of $K^{\times}$. Any representation $\rho$ affording the character $\chi$ fixes definite Hermitian forms. All these invariant forms have the same discriminant if and only if the degree $\chi(1)$ is even. Then we say that $\chi$ is unitary stable.
Let $\chi$ be such an ordinary irreducible even degree indicator $o$ character. Due to Schur indices there might not be a representation $\rho$ over the character field $L$ affording $\chi$. Nevertheless the character defines a central simple $L$-algebra with an involution of the second kind, for which [10, Section 10] defines a Brauer class $[D]$ of a quaternion algebra over $K$, the so called discriminant algebra, that allows to read off the discriminant of all $\rho(G)$-invariant Hermitian forms for any representation $\rho$ of $G$ affording the character $\chi$ (cf. Remark 4.5). If $\chi$ is quasi-split (Definition 4.4) (i.e. the quotient of degree and Schur index is even) then $L$ splits $D$ and hence $D=(L,\delta)_{K}$ for a well defined $\delta\in K^{\times}/N$, which we then call the unitary discriminant of $\chi$ (see Definitions 4.4 and 5.9). In our heavy computations we could always avoid to work with discriminant algebras; the relevant indicator $o$ characters all turned out to have Schur index $1$, i.e. they are characters of a representation over the character field. However, for the sake of future applications, we give the basics for discriminant algebras and formulate most results in these terms. Thanks to Remark 4.5 the reader may replace the discriminant algebra (which is an element of the Brauer group of $K$) by the discriminant (represented by a non-zero element of $K$).
@@
The computational methods, developed for orthogonal characters in [15], such as restriction, induction, tensor products and symmetrizations can also be used in the unitary case. Also the decomposition number techniques can be varied to obtain restrictions for the unitary discriminants (see Section 6.1), for instance showing that all primes ramifying in the discriminant algebra also do divide the group order (see Corollary 6.4).
@@
However there is a great difference to the orthogonal case: The discriminant of a given $L/K$-Hermitian form $H$ is only well defined modulo norms, $\operatorname{disc}(H)\in K^{\times}/N_{L/K}(L^{\times})$. For finite fields, the norm map is surjective, so it seems to be impossible to obtain any information about $\operatorname{disc}(H)$ by reducing the representation modulo primes. The latter technique is important in [15] to handle large degree characters, for which we computed composition factors of a condensed module (see Section 10.1) modulo some well chosen primes (usually not dividing the group order) to get enough information on the square class containing the orthogonal discriminant. It is amazing that also this condensation method can be transferred to the unitary case: If there exists a suitable automorphism $\alpha$ of order 2 of $G$, interchanging the indicator $o$ character $\chi$ with its complex conjugate $\overline{\chi}=\chi\circ\alpha$, then there is a square class of the real subfield of the character field of $\chi$, called the $\alpha$-discriminant of $\chi$, that relates to the unitary discriminant of $\chi$ (see Theorem 9.1). In complete analogy to the orthogonal case, we then can use skew adjoint units in the $\alpha$-fixed algebra to compute the $\alpha$-discriminant of $\chi$ (see Section 10.2).
@@
We start by recalling the basic notions for quadratic and Hermitian forms. As we are mainly interested in totally positive definite Hermitian forms over quadratic complex extensions $L$ of totally real number fields $K$, Section 3 gives the important facts for this situation. In particular Hasse’s norm theorem gives criteria for computing the norm subgroup $N_{L/K}(L^{\times})$ of $K^{\times}$ from local data. Invariant lattices provide a tool to use modular reduction to restrict the possible unitary discriminants. Therefore some facts about lattices in local Hermitian spaces are recalled in Section 3.2.
Section 4 recalls and extends the relevant notions for discriminants of algebras with involution. Many of the results can be found in [10, Section 10].
The next section introduces unitary discriminants and discriminant algebras of characters. To appropriately use modular reduction techniques, we need to have a finer notion of indicator for irreducible Brauer characters: There are two different sorts of indicator $o$ Brauer characters; those, for which the corresponding simple module carries a non-zero invariant Hermitian form, and those for which there is no such invariant form. The former characters are called unitary, as the corresponding representation embeds into the unitary group. To simplify notation we also say that all real Brauer characters and all ordinary characters are unitary. Then a character is called unitary stable (see Definition 5.10) if and only if all its unitary constituents have even degree. Unitary stable quasi-split ordinary characters are those that have a well defined unitary discriminant. Section 6 collects important character theoretic methods that can be automatised to compute and validate unitary discriminants. For orthogonal characters there was no need to deal with covering groups, as there the orthogonal discriminants are predicted by theoretical methods. For unitary characters the $Q_{8}$-trick described in Section 8 is a shortcut to obtain unitary discriminants of faithful characters of certain even covering groups.
@@
The next section lays the ground for applying condensation techniques to compute unitary discriminants. The discriminant of a symmetric bilinear form can be computed as the discriminant of its adjoint involution, and hence as the square class of the discriminant of any skew-adjoint unit. The adjoint involution of an Hermitian form $H$ is an involution of second kind. Also here we can obtain the discriminant of $H$ intrinsically from the algebra $\operatorname{End}(H)$ with involution (see Remark 4.5). There are two obstacles, however: The discriminant algebra from Remark 4.5 is hard to determine in general and the discriminant of $H$ is only defined modulo norms. Orthogonal subalgebras (Definition 4.8), if they exist, provide a workaround. Orthogonal subalgebras can be obtained as $\alpha$-fixed algebras (Definition 4.6) for certain outer automorphisms $\alpha$. Section 10 shows how this can be applied to compute unitary discriminants using modular condensation techniques. After reporting on highly non-trivial computations for the Harada-Norton group, we give a small example for $U_{3}(7)$ which on the one hand illustrates the condensation techniques and on the other hand generalises to compute the unitary discriminants of the groups $U_{3}(q)$ for odd prime powers $q$ (see [6]) (and very likely for other finite groups of Lie type). The last section is devoted to report on one interesting example, the group $3.ON$, the Schur cover of the sporadic simple O’Nan group. Besides giving examples for the various methods described in this paper, the results are interesting, as this is one of the rare cases, where some unitary discriminants are not rational. Non-rational orthogonal discriminants occur much more frequently (see [1]).
@@
Special thanks go to David Schlang, who compiled a large list of unitary discriminants during his master thesis [21] and his fellowship awarded by the SFB TRR 195, Computeralgebra. With his help we obtained the unitary discriminants of all indicator $o$ even degree ordinary characters for all groups in the ATLAS of finite groups [5] of order smaller or equal to the order of the Harada-Norton group, the current status of the database of orthogonal discriminants. I also thank Thomas Breuer, who is permanently updating the OSCAR database of orthogonal and unitary discriminants of characters [1].
This paper is a contribution to Project-ID 286237555 – TRR 195 – by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).
@@ 2 Quadratic and Hermitian forms
Let $L$ be a field of characteristic $\neq 2$ and $\sigma$ an automorphism of $L$ of order $1$ or $2$. Put $K:=\operatorname{Fix}_{\sigma}(L)$ to be the fixed field of $\sigma$ in $L$. Then either $\sigma=\operatorname{id}$ and $L=K$ and we talk about quadratic forms or $\sigma$ has order 2 and $[L:K]=2$. A Hermitian space, or, more precisely, an $L/K$-Hermitian space is a finite dimensional vector space $V$ over $L$ with a non-degenerate Hermitian form $H:V\times V\to L$, i.e. a $K$-bilinear map such that $H(av,w)=aH(v,w)$ and $H(v,w)=\sigma(H(w,v))$ for all $v,w\in V$ and $a\in L$. Let
@@
Of course the discriminant is already determined by the dimension and the determinant, so discriminant and determinant are equivalent invariants. However, the discriminant seems to be the more appropriate choice, mostly because discriminants of hyperbolic spaces are 1.
For quadratic forms $H$, i.e. $L=K$, the square class $\operatorname{disc}(H)=d(K^{\times})^{2}$ defines a unique étale $K$-algebra $\mathcal{D}(H)=K[X]/(X^{2}-d)$ of degree 2 over $K$, the discriminant algebra of the quadratic form (see [15, Definition 3.1, Remark 3.2] also for the correct definition for fields of characteristic $2$). For $L\neq K$ the class of the discriminant defines a unique quaternion algebra over $K$ which we call the discriminant algebra of $H$:
@@
For $a,b\in K^{\times}$ let $(a,b)_{K}$ denote the central simple $K$-algebra with $K$-basis $(1,i,j,ij)$ such that $i^{2}=a$, $j^{2}=b$, ij=jiij=-ji. If $L=K[\sqrt{\delta}]$ is a quadratic extension of $K$ we also put
@@
where $\operatorname{Br}_{2}(L,K)$ is the exponent $2$-subgroup of the Brauer group $\operatorname{Br}(K)$ of central simple $K$-algebras that are split by $L$.
@@
For $D]\in\operatorname{Br}_{2}(L,K)$ the $L$-discriminant of $[D]$ is defined as
@@
where $b\in K^{\times}$ is such that $D]=[(L,b)_{K}]$.
For an $L/K$-Hermitian form $H$ the $L$-discriminant of $\Delta(H)$ is equal to the discriminant of $H$.
@@ 3 Hermitian forms over number fields
We now assume that $L$ is a complex number field with totally real subfield $K:=L^{+}$, in particular $[L:K]=2$ and $\operatorname{id}\neq\sigma\in\operatorname{Aut}_{K}(L)$ is the non-trivial $K$-linear field automorphism of $L$. Let $(V,H)$ be a totally positive definite Hermitian space of dimension, say, $n:=\dim_{L}(V)$.
@@
The following proposition is given in [20] for general global fields. Note that in [20, Remark (10.1.4)] the determinant of the Hermitian form has to be replaced by its discriminant to obtain a correct statement.
@@
(see [20, Remark (10.1.4), Theorem (10.1.7)]) Two Hermitian spaces $(V,H)$ and $(V^{\prime},H^{\prime})$ are isometric if and only if the quadratic spaces $(V_{K},Q_{H})$ and $(V^{\prime}_{K},Q_{H^{\prime}})$ are isometric. If $L=K[\sqrt{\delta}]$ then $\operatorname{disc}(Q_{H})=\delta^{n}$ and the Clifford invariant of $Q_{H}$ is $\Delta(H)$.
@@
Two totally positive definite $L/K$ Hermitian forms are isometric, if and only if they have the same dimension and the same discriminant.
An important feature for number fields $K$ is the so called Hasse principle or local-global principle. The completions of $K$ are given by the embeddings of $K$ into the real or complex numbers (the infinite places) and the finite places, which are in bijection n to the maximal ideals of its ring of integers ${\mathbb{Z}}_{K}$.
@@
  • Two quaternion algebras over $K$ are isomorphic, if and only if they ramify at the same places of $K$. Moreover the number of ramified places is always even. Recall that a place of $K$ is ramified in the quaternion algebra $Q$, if the corresponding completion of $Q$ is a division algebra.
  • Hasse Norm Theorem for quadratic extensions: For $a\in K$ it holds that $a\in N_{L/K}(L)$ if and only if $a\in N_{L_{\wp}/K_{\wp}}(L_{\wp})$ for all places $\wp$ of $K$.
Let $\wp\unlhd{\mathbb{Z}}_{K}$ be a maximal ideal. Then there are three cases:
@@
    $\wp{\mathbb{Z}}_{L}$ is a maximal ideal of ${\mathbb{Z}}_{L}$, i.e. $\wp$ is inert in $L/K$ and $L_{\wp}$ is the unique unramified quadratic extension of $K_{\wp}$.
@@
    $\wp{\mathbb{Z}}_{L}=P_{1}P_{2}$ is a product of two distinct prime ideals $P_{1}=\sigma(P_{2})$ of ${\mathbb{Z}}_{L}$, i.e. $\wp$ is split in $L/K$ and $L_{\wp}\cong K_{\wp}\oplus K_{\wp}$.
@@
    $\wp{\mathbb{Z}}_{L}=P^{2}$ is the square of a prime ideal of ${\mathbb{Z}}_{L}$. Then $\wp$ is ramified in $L/K$.
@@
    $\nu_{\wp}(a)$ is even for all prime ideals $\wp$ of $K$ that are inert in $L/K$
@@
The first criterion guarantees that $a$ is a norm for all infinite places of $K$, the second one that $a$ is a local norm for all inert places and the third one for the ramified places. For the split places $L_{\wp}=K_{\wp}\oplus K_{\wp}$ and all elements of $K_{\wp}$ are norms. ∎
For the real completions the signature parameterizes Hermitian forms. As $H$ is totally positive definite, this specifies $H$ for the real completions of $K$.
@@
The infinite places of $K$ are ramified in the discriminant algebra $\Delta(H)$ if and only if $\operatorname{disc}(H)$ consists of totally negative elements, i.e. $\dim_{L}(V)\equiv$ 22 or $3$ mod $4$.
Let $\wp$ be a prime ideal of $K$ that is split in $L/K$. Then the completion $\Delta(H)_{\wp}$ contains the completion $L_{\wp}\cong K_{\wp}\oplus K_{\wp}$ which contains zero divisors. So $\wp$ is not ramified in $\Delta(H)$.
@@
Let $\wp$ be a prime ideal of $K$ that is inert in $L/K$. Then $\wp$ is ramified in $\Delta(H)$ if and only if the $\wp$-adic valuation $\nu_{\wp}(d)$ is odd for all $d\in\operatorname{disc}(H)$.
@@
Now $\wp$ is ramified in $\Delta(H)=[(L,d)_{K}]$ if and only if $d$ is not a norm in $L_{\wp}/K_{\wp}$, i.e. if and only if $\nu_{\wp}(d)$ is odd. ∎
@@
For a ramified place $\wp=P^{2}$ we can always find a representative $d\in{\mathbb{Z}}_{K}$ of $\operatorname{disc}(H)$ such that $d\not\in\wp$ (see Proposition 3.10 (c)). Assume that $\wp$ is not a dyadic prime, i.e. $2\not\in\wp$. Then $\wp$ ramifies in the discriminant algebra of $H$ if and only if $d$ is not a square modulo $\wp$.